OpenStudy (anonymous):

please check the points for the equation x^2/8 + y^2/5 = 1 (-8,0),(8,0),(0,-5),(0,5)

ganeshie8 (ganeshie8):

i think squares should be there for the bottom numbers too ?

ganeshie8 (ganeshie8):

\[\large \dfrac{x^2}{8} + \dfrac{y^2}{5}=1\] or \[\large \dfrac{x^2}{8^2} + \dfrac{y^2}{5^2}=1\]

ganeshie8 (ganeshie8):

which one is correct ?

OpenStudy (anonymous):

x^2/8 + y^2/5 = 1

ganeshie8 (ganeshie8):

then none of the given points satisfy the equation plug them and see

ganeshie8 (ganeshie8):

x^2/8 + y^2/5 = 1 (-8,0) plugin x = -8, y = 0 (-8)^2/8 + 0^2/5 = 1 8 = 1

ganeshie8 (ganeshie8):

you need to provide more details about the question

OpenStudy (anonymous):

all it gives is the points

ganeshie8 (ganeshie8):

can u take a screenshot and attach if psble?

OpenStudy (anonymous):

so it would have to be x^2/64 + y^2/25=1

ganeshie8 (ganeshie8):

i need a screenshot

ganeshie8 (ganeshie8):

thnks! now it makes sense

ganeshie8 (ganeshie8):

yes D is right !